OpenStudy (chupacabraj):

Can someone please check my work? <3

OpenStudy (chupacabraj):

Problem: Pennies today are copper-covered zinc, but older pennies re 3.1 g of solid copper. What are the total positive charge and total negative charge in a solid copper penny that is electrically neutral?

OpenStudy (chupacabraj):

My attempt at solving it: \[3.1 g Cu \times \frac{ 1 mol Cu }{ 63.54 g Cu }\times \frac{ 6.022 x 10^{23} }{ 1 mol Cu }= 2.9 \times10^22 atoms of Cu\]

OpenStudy (chupacabraj):

then multiply the # of atoms of Cu by 29 (# of protons and electrons) = 8.41 e^23 that times 1.6 x 10^19 Coulomb

OpenStudy (chupacabraj):

I get an answer of 134560 Coulombs?